### v1

# Universal Debiased Editing for Fair Medical Image Classification

Abstract
In the era of Foundation Models’ (FMs) rising prominence in AI, our study addresses the challenge of biases in medical images while using FM API, particularly spurious correlations between pixels and sensitive attributes. Traditional methods for bias mitigation face limitations due to the restricted access to web-hosted FMs and difficulties in addressing the underlying bias encoded within the FM API. We propose an U(niversal) D(ebiased) E(diting) strategy, termed UDE, which generates UDE noise to mask such spurious correlation. UDE is capable of mitigating bias both within the FM API embedding and the images themselves. Furthermore, UDE is suitable for both white-box and black-box FM APIs, where we introduced G(reedy) (Z)eroth-O(rder) (GeZO) optimization for it when the gradient is inaccessible in black-box APIs. Our whole pipeline enables fairness-aware image editing that can be applied across various medical contexts without requiring direct model manipulation or significant computational resources. Our empirical results demonstrate the method’s effectiveness in maintaining fairness and utility across different patient groups and diseases. In the era of AI-driven medicine, this work contributes to making healthcare diagnostics more equitable, showcasing a practical solution for bias mitigation in pre-trained image FMs.
footnotetext: This work is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), Compute Canada and Vector Institute.
Corresponding author: Xiaoxiao Li (xiaoxiao.li@ece.ubc.ca)
1 Introduction
Using pre-trained models or encoders to convert complex input data into vector representations is widely used in computer vision and natural language processing (NLP) [26, 7]. This process transforms the original data into a representative low-dimensional hidden space, preserving information for downstream tasks. With the advancement of Foundation Models (FM), services like Google MedLM [18], Voyage.ai, and ChatGPT provide data embedding services, eliminating the need for specialized hardware or extensive training. Their effectiveness is especially notable in medical fields, addressing data scarcity and hardware constraints. Clinics can enhance these models with minimal effort by fine-tuning custom classifiers on compact embedding output from FM application programming interface (APIs), which provides machine learning (ML) service as a function without users to train their own models from scratch. However, the inherent biases in the APIs’ training data and their model potentially harm marginalized groups by perpetuating gender, race, and other biases, a problem that has been exposed in the NLP field [17]. The biased embedding can compromise models’ performance on minority groups [4]. Therefore, it is essential to develop new solutions to attain fairness in medical image embedding from pre-trained FM API.
Studies have been done to address the fairness issues in classification problems using machine learning (ML) models. These methods can be categorized into three approaches: 1)Model-based strategies update or remove bias-related model parameters to mitigate bias, for example, using adversarial methods [1, 22, 10]; or prune parameters that are significant for sensitive attributes (SAs) [23]; or update model to minimize the mutual information between target and SA representations using disentanglement learning [3]. However, the pretrained FM API services offer users very limited control over the model parameters [8, 9]. Thus model-based approaches are infeasible under the constraints. 2)Prediction calibration-based methods analyze the prediction probability distribution of a classifier and apply different thresholds to each subgroup to reduce the discrepancy among subgroups [5, 14, 16]. These methods need custom thresholds for various tasks, classes, and groups, which limits their generalizability. Moreover, a significant amount of validation data is needed to determine the thresholds, rendering them less efficient in medical imaging applications where data scarcity is an unavoidable problem. 3)Data-based strategies alleviate bias at the pre-training stage. Re-distribution and re-weighting methods [15, 16] address unfairness by adjusting the balance of subgroups. The effectiveness of redistributed training data is limited because it can only affect the classification head, not the pre-trained FM API embedding encoder. Recent data editing methods show strong ability to reduce disparity among subgroups by removing sensitive information from the input images [10, 21]. However, can leave the model unchanged, they fail to address the underlying bias encoded within the FM API. Moreover, these approaches either depend on disease labels and require significant computational costs. Yao et al. introduce an image editing method independent on targeted downstream tasks via sketching [25], but this method suffers from subpar performance as not learnable.
In this work, we take two unique properties of pretrained FM API into consideration. Firstly, we aim to emphasize the need to maintain the flexibility of using FM without limiting it to a specific task while maintaining good utility, thus focusing on the learnable data editing-based method that is independent of the downstream task. We choose to edit on the image space rather than on the FM API embeddings since the former provides better control on medical image fidelity and enables interpretability on the applied editing (as shown in Sec. 3.1). Secondly, given the constraint against modifying parameters in the pre-trained API, we commit to not changing the FM API model parameters, even with black-box access to the API. To the best of our knowledge, no effective and universal [^1] We refer ‘universal’ as debiasing API embeddings for various classification tasks. method has been available to address this important, timely, but challenging research question: How to remove bias on medical images when using their embeddings from a pre-trained FM API for various classification tasks?
To answer this question, we propose the Universal Debiased Editing, UDE, where the resulted edit can be shared across all subjects to eliminate SA-related information for various disease classification tasks. Specifically, we start by using a pre-trained SA classifier, trained with the same FM API embeddings. Then, we optimize a set of learnable parameters (referred UDE noise) to be added to the images by confusing the SA classifier. Furthermore, we introduce a greedy zeroth-order optimization strategy, GeZO, when APIs restrict gradient propagation in the black-box setting. Lastly, we apply this UDE noise to input images to generate fair embeddings and achieve unbiased disease classification across various disease tasks. Extensive evaluations on disease classification tasks show our method’s effectiveness in promoting fairness while preserving utility.
2 Method
2.1 Problem Setting
Figure 1: Overview of universal debias editing pipeline. (a) We eliminate the spurious correlation by breaking the connection between SA $\mathcal{A}$ and image $\mathcal{X}$, ensuring the model relies solely on disease-related information $\mathcal{Y}$. (b) Training UDE noise deceives the pre-trained SA classifier in $\mathcal{A}$ using a frozen FM API and SA classifier, suitable for both white-box and black-box API scenarios based on gradient accessibility. (c) We demonstrate the use of UDE noise: users augment their images with this noise, extract embeddings via the FM API, and proceed to train fair disease classifiers. (d) The (G)reedy (Z)eroth-(O)rder (GeZO) black-box editing method, selects the optimal perturbation via the Greedy Gradient process when gradients is inaccessible. It tracks the best perturbation using velocity in each local iteration to update the UDE noise $\epsilon$.
This section outlines the problem of fairness in a binary medical image classification task. Firstly, we define key variables: input images $x\in\mathcal{X}$, binary disease labels $y\in\mathcal{Y}=\{0,1\}$, and sensitive attribute $a\in\mathcal{A}=\{1,...,|\mathcal{A}|\}$ (i.e. $|\mathcal{A}|=2$ for gender with male and female). As patients may not have sensitive and disease labels simultaneously, we denote images with sensitive attribute labels as $\{x_{i},a_{i}\}^{N}=[X_{A},A]$ and images with disease label as $\{x_{j},y_{j}\}^{M}=[X_{T},Y]$, where $N$ and $M$ are the number of samples. Then we denote the FM API as $\phi:\mathcal{X}\rightarrow\mathcal{Z}$, to obtain the image embedding $z=\phi(x)$; the disease classifier as $f:\mathcal{Z}\rightarrow\mathcal{Y}$ and SA classifier as $g:\mathcal{Z}\rightarrow\mathcal{A}$.
Fairness Issue: Fig. 1 (a) shows there exists an association between sensitive attributes $\mathcal{A}$ and targets $\mathcal{Y}$ (dashed line). In the Empirical Risk Minimization (ERM) [20], the model $f(\phi(\cdot))$ superficially consider $\mathcal{A}$ as a proxy of $\mathcal{Y}$, causing the issue of fairness. This phenomenon persists even when there is an equal number of data points for each class in both $\mathcal{Y}$ and $\mathcal{A}$, as demonstrated in Sec. 3.1. Our approach aims to address these disparities using universal debiased edit.
2.2 Universal Debias Editing on Image for Fair Disease Classification
In this section, we present our image editing approach aimed at reducing bias via learning a universal noise tensor with the same dimensionality as the image, which is called UDE noise. Then, we explain how to use this UDE noise for fair disease classification.
Universal Debiased Editing. Given the spurious correlation between SA $\mathcal{A}$ and disease $\mathcal{Y}$, we aim to remove this link through intervention on $\mathcal{X}$ by adding minimal universal editing $\epsilon$, known as UDE noise, concealing the spurious correlations while preserving image utility. As shown in Fig. 1 (b), we first obtain a pre-trained SA classifier for medical images by either leveraging an existing one (if available) or using the collected $[X_{A},A]$ to train the classifier $g$ with cross-entropy loss. Then, we freeze the classifier and leverage gradient ascent to learn the minimal UDE noise $\epsilon$, which is constrained by certain threshold to preserve image fidelity. $\epsilon$ is trained to deceive the SA classifier $g$ without compromising useful information by solving the following optimization problem:
$\mathcal{L}_{\epsilon}=-\left[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\texttt{CE}}(a_{i},g(\phi(x_{i}+\epsilon)))\right]+\lambda||\epsilon||_{2}$ (1)
where $\lambda$ regularizes the magnitude of $\epsilon$. Sec. 3.4 explores its effect in detail. Importantly, we also offer a zero-order strategy for black-box APIs, as detailed in Sec. 2.3.
Fair Disease Classification Having obtained the UDE noise $\epsilon$ to conceal the SA $\mathcal{A}$, we then freeze and add it to disease classification images $X_{T}$. As shown in Fig. 1 (c), this action guarantees that the FM generates a fair embedding devoid of sensitive information, denoted as $\hat{z}_{i}=\phi(x_{i}+\epsilon)$. Finally, these fair embeddings are leveraged by the subsequent disease classifier $f$ for prediction. To train the classifier $f$, we employ a cross-entropy loss:
$\mathcal{L}_{T}=\frac{1}{M}\sum_{i=1}^{M}\mathcal{L}_{\texttt{CE}}(y_{i},f(\phi(x_{i}+\epsilon)),$ (2)
where classifier $f$ is the only trainable parameter and $M$ is the size of $X_{T}$.
2.3 Greedy Zero-order Optimization for Black-box FM API
Zero-order (ZO) optimizations [2, 12], which avoid the need for gradient computations, offer distinct advantages in optimizing black-box models. However, these methods often exhibit slower convergence rates, particularly in training large FM models. In contrast, our proposed UDE, concentrating on the input space, offers a mitigation of this training efficiency issue of ZO optimization.
Inspired by ZO-SGD [19] and recent MeZO [13] that employ in-place perturbation updates, we propose a G(reedy) Z(ero-)O(rder) (GeZO) optimization specifically for efficient UDE. GeZO employs in-place perturbations [19, 13] and greedily updates with the gradient sign that achieves global optimal loss, thereby accelerating the optimization process. As shown in Fig. 1 (d), the core procedure of GeZO is the Greedy Gradient in the right box, where the gradient of UDE’s objective (Eq.  1) is estimated for each local iteration $r\in\{1,...,|R|\}$ in GeZO. Greedy Gradient takes the UDE noise from the current local iteration ($\epsilon^{r}$) and estimates its gradient by continuously adding minor perturbations ($\delta$) in different directions (e.g., $d_{1}$ and $d_{2}$) to it, i.e., $d_{1}\cdot\delta$ and $d_{2}\cdot\delta$. It then greedily selects the best direction perturbation ($d_{\texttt{best}}\cdot\delta$) that results in the smallest objective in this process. Like MeZO [13], we integrate stochastic sampling for $C$ times. Each time we calculate the objective using a subset of training data to avoid sucking to the local optimum. Once the best direction and perturbation are selected for each local iteration $r$, we add $d_{\texttt{best}}\cdot\delta$ to the current velocity, $v^{r}$. This velocity keeps track of the accumulated updating direction and magnitude for $\epsilon$ for each local iteration $r$, where the initial value for velocity $v^{1}$ is 0. As shown in the right part of Fig. 1 (d), $v^{r+1}$ is updated by adding the best perturbation returned by the Greedy Gradient. Finally, the UDE noise $\epsilon^{r+1}$ is updated through adding the $v^{r+1}$. The local iterations $|R|$ is a hyper-parameter, where a smaller iteration increases the efficiency at the cost of the more biased gradient estimation. We investigate the effect of different $|R|$ in Sec. 3.4. In actual implementation, we also introduced a momentum to update the velocity to accelerate the convergence. While the key idea is presented in this section, we provide a detailed step-by-step algorithm box for GeZO in Alg. 1.
Algorithm 1 Greedy Zero-order (GeZO) Optimization
Setup: Local Iterations $R$, step size $s$, step size decay $k$, Edit from last global epoch $\epsilon_{\texttt{epoch}}$, sample times $C$, best loss $\mathcal{L}_{\texttt{best}}$, best direction $d_{\texttt{best}}$, momentum: $\mu$, batch size $B$.
1:procedure UpdateEditionPerEpoch($\epsilon_{\texttt{epoch}}$):
2:     $\epsilon^{1}\leftarrow\epsilon_{\texttt{epoch}}$, $s^{1}\leftarrow 0.01$, $v^{1}\leftarrow 0$, $\mu\leftarrow 0.9$, $\mathcal{L}_{\texttt{best}}\leftarrow\infty$, $k\leftarrow 0.95$
3:     for $r=1\to R$ do
4:         $d_{\texttt{best}},\mathcal{L}_{\texttt{best}}\leftarrow$GreedyGradient($B,\epsilon^{r},s^{r}$) $\triangleright$ find the $d_{\texttt{best}}$ achieve $\mathcal{L}_{\texttt{best}}$
5:         if $d_{\texttt{best}}\neq\text{None}$ then
6:              $v^{r+1}\leftarrow\mu\cdot v^{r}+d_{\texttt{best}}$
7:              $\epsilon^{r+1}\leftarrow\epsilon^{r}+v^{r+1}$, $s^{r+1}\leftarrow s^{r}$triangleright Update $\epsilon$ with momentum
8:         else
9:              $s^{r+1}\leftarrow k\cdot s^{r}$triangleright Reduce step size if no improvement
10:     return $\epsilon^{R}$triangleright Return the updated $\epsilon_{\texttt{epoch + 1}}$
11:procedure GreedyGradient($B,\epsilon^{r},s^{r}$)
12:     $d_{\texttt{best}}\leftarrow\text{None}$
13:     for $c=1\to C$ do $\triangleright$ Sample $C$ times
14:         $\delta\leftarrow\text{RandomPerturbation}()\cdot s_{r}$triangleright Generate scaled perturbation
15:         for $d\in\{-1,1\}$ do
16:              $\epsilon^{\prime}\leftarrow\epsilon^{r}+d\cdot\delta$triangleright Apply perturbation in both directions
17:              $\mathcal{L}_{\epsilon}=-\left[\frac{1}{B}\sum_{i=1}^{B}\mathcal{L}_{\texttt{CE}}(a_{i},h(\phi(x_{i}+\epsilon^{\prime})))\right]+\lambda||\epsilon^{\prime}||_{2}$triangleright One batch
18:              if $\mathcal{L}_{\epsilon}<\mathcal{L}_{\texttt{best}}$ then
19:                  $\mathcal{L}_{\texttt{best}}\leftarrow\mathcal{L}_{\epsilon}$, $d_{\texttt{best}}\leftarrow d\cdot\delta$triangleright Update best loss and direction
20:     return $d_{\texttt{best}},\mathcal{L}_{\texttt{best}}$.
3 Experiement
3.1 Settings
Table 1: Training and Testing Set Distribution for Pneumonia, Edema, and Pleural Effusion.
Diseases Pneumonia Edema Pleural Effusion
Negative Positive Negative Positive Negative Positive
M F M F M F M F M F M F
##\## Train Sample 1500 150 150 1500 5000 500 500 5000 5000 500 500 5000
##\## Test Sample 100 100 100 100 200 200 200 200 200 200 200 200
Dataset. To demonstrate the generalizability of UDE. we adopt the CheXpert dataset [6], a chest X-ray dataset with multiple disease labels, to predict the binary label for Pleural Effusion, Pneumonia and Edema individually in chest radiographs. All the ablation studies are performed on Pleural Effusion classification task for space limit. We take gender bias (male and female) as an example due to its broad impact on society and medical imaging analysis. To demonstrate the effectiveness of bias mitigation methods, we follow [3] to amplify the training data bias for each disease by (1) firstly dividing the data into different groups according to the SA; (2) secondly calculating the positive rate of each subgroup; (3) sampling subsets from the original training dataset and increase each subgroup’s bias gap (more positive sample in a subgroup). Then, we sample testing data with the same procedure, but achieve an equal subgroup bias gap.
Evaluation metrics. We use the classification accuracy to evaluate the utility of classifiers on the test set. To measure fairness, we employ equal opportunity (EO) [5] and disparate impact (DI) [11] metrics. EO aims to ensure equitable prediction probabilities across different groups, defined by the sensitive attribute $a^{j}$, for a given class $y^{j}$. It quantifies the disparity in true positive rates between groups: $\text{EO}_{Y=y_{1}}=P(\hat{Y}=y^{1}|Y=y^{1},A=a^{1})-P(\hat{Y}=y^{1}|Y=y^{1},A=a^{2})$, where a smaller gap signifies greater equality of opportunity. DI evaluates the presence of indirect discrimination by measuring the ratio of positive predictions across different groups: $\text{DI}=\frac{P(\hat{Y}=1|A=a^{1})}{P(\hat{Y}=1|A=a^{2})}$. DI closing to 1 indicates the minimal disparity in positive prediction rates between the groups. To align with EO metric, we employ the $|1-\textsc{DI}|$ to quantify the fairness in percentage, with smaller values denoting greater fairness.
3.2 Implementation Details
Architecture. In our implementation, all methods use the same architecture. To simulate the FM API, we utilize a publicly available self-supervised pre-trained ViT-base model  [^2] https://github.com/lambert-x/Medical_MAE, optimized on X-ray images, given its superior performance metrics among all other architectures [24]. Within our study, we fix the encoder of the ViT model and only fine-tune the classification head, simulating the configuration detailed in Sec. 2.1.
Univeresal Debiased Editing. For UDE, we first fine-tune the SA classifier, $g$, using the feature encoded by the FM ($\phi(x_{i})$) with the training set of Pleural Effusion in Table 1. We update it using Adam with a learning rate (lr) of $0.0001$ for $50$ epochs. Second, we update the UDE noise through implementing Eq. 1 by initializing $\epsilon$ as a trainable PyTorch parameter with all entries initially set to zero. The parameter matches the input image dimensions of $224\times 224$. If not otherwise mentioned, we choose $\lambda=0.01$ in Eq. 1. It is optimized using the Adam optimizer with lr of $0.01$ for $50$ epochs. Finally, we fine-tune the disease classifier, $f$, with $\epsilon$ added on the input data as delineated in Eq. 2. We update it using AdamW with lr of $1.25\times 10^{-4}$ for $50$ epochs.
GeZO. Here, we keep all parts for UDE the same as above, except using GeZO to update the edit rather than the Adam, given that the gradient is not accessible. The only key parameter that we are interested in is the local iteration $T$, where we set $T=10$ if not otherwise mentioned. Furthermore, the detailed implementation of the algorithm and hyperparameter setting for GeZO can be found in Alg. 1.
3.3 Comparison with Baselines
Table 2: Comparision of binary prediction of Pleural Effusion, Pneumonia, and Edema. We label the best performance in bold and the second-best performance with underline. All the values are in percentage.
Diseases Pleural Effusion Pneumonia Edema
$EO_{n}\downarrow$ $EO_{p}\downarrow$ DI $\downarrow$ Acc $\uparrow$ $EO_{n}\downarrow$ $EO_{p}\downarrow$ DI $\downarrow$ Acc $\uparrow$ $EO_{n}\downarrow$ $EO_{p}\downarrow$ DI $\downarrow$ Acc $\uparrow$
ERM 40.5 57.0 58.5 72.9 70.0 70.0 74.5 59.6 42.0 39.0 42.0 74.5
Sketch [25] 44.0 52.0 57.5 66.3 64.0 61.0 72.6 56.3 44.0 15.0 15.5 67.0
Group DRO [16] 41.0 58.0 59.5 72.3 60.0 56.0 64.4 61.0 40.5 40.5 43.3 74.8
Batch Samp. [15] 41.5 50.5 51.8 74.3 64.0 72.0 76.6 60.5 39.0 43.0 44.8 76.0
BiasAdv [10] 39.0 54.5 58.2 71.1 36.0 54.0 77.1 61.0 39.0 33.5 36.6 74.9
UDE 28.5 23.0 25.6 75.1 38.0 25.0 34.7 61.3 27.5 17.0 19.7 75.6
UDE-GeZO 37.0 25.0 28.5 72.8 35.0 27.0 33.7 61.5 34.0 15.5 17.2 75.1
3.3.1 Quantitative Analysis.
Table 2 summarizes the performances among different diseases and baselines. The baselines include model-based strategy, like biasAdv  [^3] In BiasAdv’s implementation, we treat FM API as a white box as it requires access to FM’s gradient and there is no zeroth-order optimization for it yet. [10]; data-based strategies, like batch sampling [15] with data re-distribution and sketch [25] with data generation; and prediction calibration-based strategy, like Group DRO [16]. The details of these baselines are introduced in Sec. 1. The results indicate that the UDE effectively balances fairness and utility. Taking Pleural Effusion as the example to analyze, the white-box optimized UDE outperforms all baselines. The $EO_{p}$ and DI decrease over 25% compared to all the baselines, an indication of less disparity between the two genders. Additionally, the utility not only maintains but also surpasses all other groups, e.g., UDE’s accuracy (Acc) is higher than 2.2% compared to ERM. Given we are using a balanced testing set, this increase is also a sign of better generalization. Similar for Pneumonia and Edema, where UDE and UDE-GeZO takes the best two performance for almost all entries. Although learned through black-box optimization, our UDE-GeZO achieves comparable performance to the standard optimization used in UDE. This demonstrates not only the validity of GeZO but also the efficiency afforded by the small number of edition parameters [13]. In Pneumonia, UDE-GeZO even slightly surpasses the performance of UDE across most metrics.
These findings affirm the efficacy of UDE, where it not only facilitates fair medical image embedding and training but also introduces better generalizability in downstream classification tasks. Furthermore, the universal applicability of the UDE is validated in different diseases, where it consistently outperforms the baselines across three diseases in both fairness and utility.
3.4 Ablation Studies
Effect of Regularization Coefficient. To investigate the effect of different regularization coefficients’ ($\lambda$) effect, we vary $\lambda$ from 0 to 1. Fig. 2 (a) depicts the EO and the Acc for different $\lambda$s. The metrics for ERM are labeled as horizontal dashed lines for convenience. As shown, both the fairness and utility metrics remain relatively stable for $\lambda<1$, consistently surpassing the ERM baseline. However, as $\lambda$ increases to 1, we observe a marked decline in accuracy below the ERM benchmark, along with a notable increase in the EO metric for both classes. This performance deterioration can be explained by the excessive constraint imposed by the larger $\lambda$, which severely limits the magnitude of editing, thus affecting the UDE’s functionality to learn useful disease features to improve fairness and accuracy.
Refer to caption (a)
Refer to caption (b)
Figure 2: Ablation study of our methods: (a) Effect of different $\lambda$; (b) Effect of different local epochs using GeZO.
Effect of Local Iterations in GeZO. In black-box API, total local iterations $R$ in GeZO affect the optimization performance, wherein larger $R$ leads to more accurate optimization at the cost of efficiency. Here, we examine the effect of changing $R$, as shown in Fig. 2 (b). For fairness metrics, increasing $R$ from 2 to 20 significantly reduces the EO score for both classes, demonstrating a considerable debiasing impact with more local epochs. This effect is attributed to increased perturbation sampling that expands the search space with more local epochs, as introduced in Sec. 2.3. Meanwhile, accuracy remains relatively stable, with minor fluctuations between 72% and 73%. Based on these findings, we choose $R=10$ for an optimal balance between fairness and performance.
3.5 Visualization on UDE Noise
(a) (a)
(b) (b)
(c) (c)
Figure 3: Visualization of images with UDE noise with gaussian smoothing for (a) Pleural Effusion; (b) Pneumonia; (c) Edema. For each row, images are sequenced from left to right as follows: negative and positive images optimized via white-box when the gradient is accessible, followed by negative and positive images optimized via GeZO when gradient access is constrained. We note that adding UDE noise onto the original image still provides visually high fidelity.
Figure 4: Visualization of chest X-rays and UDE noise patterns (with gaussian smoothing) to interpret gender-discriminative image regions. (a) The normalized UDE noise map, with larger noise highlighted by brighter color, reveals gender-discriminative features. The large noise circled in orange corresponds to the breast. The large noise circled in yellow reflects artifacts on X-ray, such as text notations. (b) A female chest X-ray. (c) A male chest X-ray.
This section presents the visualization of our added noise and images. Fig. 3 illustrates the preservation of high fidelity in images after adding noise, while Fig. 4 interprets this noise. To enhance visual clarity, we apply a Gaussian kernel to smooth out the noise, addressing the patch-like patterns that arise from using a ViT as the encoder.
Fig. 3 demonstrates that images modified with noise maintain good fidelity for various diseases, under both white-box and black-box optimization methods. This suggests that the modified images continue to convey critical visual disease information for human evaluation.
Fig. 4 (a) displays the top 20% of the normalized noise map, highlighting larger noises in brighter colors and marking areas with significant noise magnitude for easy identification. Fig. 4 (b) and (c) examine chest X-rays of a female and a male, respectively. Notably, large noises, marked in orange, align with areas potentially containing gender-specific information, such as the breast region. Yellow circles highlight areas of artifacts, like machine annotations on the X-ray. These noise maps enhance the interpretability of SA discriminative features by clearly identifying significant noise locations.
4 Conclusion
In this study, we address a crucial, yet under-explored aspect of health equity—the inherent bias in FM API’s usage for classification—through the introduction of universal debiased editing. UDE effectively masks bias-inducing pixels, enhancing fairness in API-generated embeddings. Furthermore, GeZO tackles the challenge of the inaccessibility of the gradient in black-box APIs by estimating gradients via perturbation. Future research will extend UDE’s application across various FM APIs and settings, aiming to solidify its role in promoting fairer machine learning practices.